# 2cm2.5cm2.5cm2cm

### v2

@@ 6 Conclusions
The upper bound for the term given in (A-1) follows exactly as discussed in Liu (2012): choose $\epsilon^{\prime}>0$ and $\eta>0$ small enough such as $\epsilon^{\prime}+\dfrac{8\eta}{1-\|{\bf A}\|_{p}}<\epsilon$ and $\|{\bf x}-{\bf z}\|_{p}<\eta$ implying $|w({\bf x})-w({\bf z})|<\epsilon^{\prime}$, with $p\in[1,\infty]$. In this part, we are using the fact that $\|{\bf A}\|_{p}<1$ for some $p\in[1,\infty]$, which follows from the assumption that there exists $p$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$.
By combining the above results, we obtain that $\left|E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf x})-E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf z})\right|<\epsilon^{\prime}+2\|{\bf x}-{\bf z}\|_{p}$. For general $k\geq 2$, the result follows by using mathematical induction, exactly as done in Chapter 4 from Liu (2012), and therefore it is omitted. With the e-chain property established for the bivariate process $\{\boldsymbol{\xi}_{t}\}$ and under the assumption of existence of $p\in[1,\infty]$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, following the same steps that proofs of Proposition 4.2.1(b) and Proposition 4.3.1 from Liu (2012), we obtain the desired result.
@@
Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a first-order negative binomial INGARCH process. Under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.1 holds (for $p=1$) and gives us that $\{Y_{t}\}_{t\in\mathbb{N}}$ has a unique stationary and ergodic solution. Moreover, this process is a Markov chain. Therefore, we will now show that Tweedie’s criterion is satisfied for a test function to show the finiteness of arbitrary moments stated in the lemma (Meyn and Tweedie, 1993).
@@
with $E(Y)=\lambda$.
From (A-5), we immediately obtain that $E(Y^{d})=\lambda^{d}(1+1/\phi)+O(\lambda^{d-1})$ as $\lambda\rightarrow\infty$. Using this result, it follows that
@@
Hence, we can express the conditional expectation (A-4) as
@@
as $r\rightarrow\infty$, where the last inequality follows from lemma’s assumption implying that $\beta_{1}+\beta_{2}<1$ and $\alpha_{1}+\alpha_{2}<1$. Further, the remaining term converges to 0 as $r\rightarrow\infty$: $\dfrac{O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}=O(r^{k-1})/O(r^{k})=O(r^{-1})$.
Therefore, there exist real constants $\kappa_{1}\in(0,1)$, $\kappa_{2}>0$, and $L>0$ such that (A-4) is bounded above as follows:
@@
In other words, Tweedie’s criterion is satisfied and we conclude that the $k$-th moment of $\lambda_{t}$ and $\phi_{t}$ are finite for arbitrary $k\in\mathbb{N}$. Now, using the fact that $Y_{t}|\mathcal{F}_{t-1}\sim\mbox{NB}(\lambda_{t},\phi_{t})$ and Eq. (A-5), we have that the $k$-th conditional moment of $Y_{t}$ given $\mathcal{F}_{t-1}$ is a polynomial on $\lambda_{t}$ and $1/\phi_{t}$; note that $1/\phi_{t}\leq 1/\alpha_{0}$ for all $t$. Since the $l$-th moment of $\lambda_{t}$ is finite for arbitrary $l\in\mathbb{N}$, we obtain that the unconditional $k$-th moment of $Y_{t}$ is finite for all $k\geq 1$ as well.
@@
We will show that the result holds for the third derivative of $\ell(\cdot)$ with respect to $\beta_{1}$ and also $\alpha_{1}$. The remaining cases follow the very same steps as these two and therefore are omitted.
@@
The case for $\alpha_{1}$ is more involving. We have that $\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\alpha_{1}^{3}}=\dfrac{\partial^{2}S_{2,t}}{\partial\phi_{t}^{2}}\left(\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right)^{3}$, where
@@
Inequality (15) from Guo and Qi (2013) gives that $|\Psi^{\prime\prime}(x)|<\dfrac{1}{x}+\dfrac{2}{x^{3}}$, for $x>0$. Hence, it follows that $|\Psi^{\prime\prime}(y_{t}+\phi_{t})|<\dfrac{1}{(y_{t}+\phi_{t})^{2}}+\dfrac{2}{(y_{t}+\phi_{t})^{3}}\leq\dfrac{1}{\phi_{t}^{2}}+\dfrac{2}{\phi_{t}^{3}}\leq\dfrac{1}{\alpha_{0}^{2}}+\dfrac{2}{\alpha_{0}^{3}}\leq\dfrac{1}{\alpha_{0,inf}^{2}}+\dfrac{2}{\alpha_{0,inf}^{3}}$. Similarly, $|\Psi^{\prime\prime}(\phi_{t})|<\dfrac{1}{\alpha_{0,inf}^{2}}+\dfrac{2}{\alpha_{0,inf}^{3}}$. The modulus of the remaining terms in (A-7) are bounded above (as done for the third derivative involving $\beta_{1}$) by a linear function of $y_{t}$ with positive coefficients do not depend on $\boldsymbol{\theta}$. Moreover, $\left|\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right|\leq\displaystyle\sum_{j=1}^{t-1}\alpha_{2,up}^{j-1}y_{t-j}\equiv\mu_{2t}$.
@@
A-3 Proof of Theorem 3.1
We will show that the Conditions (A1), (A2), and (A3) from Lemma 3.1 by Jensen and Rahbek (2004) are satisfied, which give us the desired results stated in the theorem. As argued before for the general case, we use Proposition 3.1 and the Central Limit Theorem for martingale difference to establish the weak convergence involving the score function in (5). The existence and finiteness of the matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ follow from Theorem 2.2 that provides the finiteness of the moments of all orders for the linear NB tv-DINARCH process under the assumption that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, which is in force. Therefore, Condition (A1) holds.
Moreover, under the assumption $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, we obtain from Theorem 2.1 that $\{Y_{t}\}_{t\in\mathbb{N}}$ is a stationary and ergodic sequence. Then, the Law of Large Numbers for stationary and ergodic sequences gives us that (6) is valid, with the existence and finiteness of the matrix $\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})$ being ensured again by Theorem 2.2. That is, Condition (A2) from Lemma 3.1 by Jensen and Rahbek (2004) is satisfied. Finally, Condition (A3) has been established in Lemma A-1.

### v3

@@ 6 Conclusions
The upper bound for the term given in (A-1) follows exactly as discussed in Liu (2012): choose $\epsilon^{\prime}>0$ and $\eta>0$ small enough such as $\epsilon^{\prime}+\dfrac{8\eta}{1-\|{\bf A}\|_{p}}<\epsilon$ and $\|{\bf x}-{\bf z}\|_{p}<\eta$ implies $|w({\bf x})-w({\bf z})|<\epsilon^{\prime}$, with $p\in[1,\infty]$. In this part, we are using that $\|{\bf A}\|_{p}<1$ for some $p\in[1,\infty]$.
By combining the above results, we obtain that $\left|E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf x})-E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf z})\right|<\epsilon^{\prime}+2\|{\bf x}-{\bf z}\|_{p}$. For general $k\geq 2$, the result follows by using mathematical induction, exactly as done in Chapter 4 of Liu (2012), and therefore it is omitted. With the e-chain property established for the bivariate process $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$, we show part (b) by following similar arguments of the proofs of Proposition 4.2.1(b) and Proposition 4.3.1 of Liu (2012) to establish the geometric moment contraction (GMC) of $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$.
Denote by $(Y_{t},\boldsymbol{\xi}_{t})$ and $(Y_{t}^{(0)},\boldsymbol{\xi}_{t}^{(0)})$ the process under the initial value $\boldsymbol{\xi}_{0}=\boldsymbol{\xi}=(\lambda,\phi)^{\top}$ and $\boldsymbol{\xi}_{0}=\boldsymbol{\xi}^{(0)}=(\lambda^{(0)},\phi^{(0)})^{\top}$, respectively. It follows that
$E\left(\|\boldsymbol{\xi}_{1}-\boldsymbol{\xi}_{1}^{(0)}\|_{p}\right)$ $=$ $E\left(\|(\boldsymbol{\tau}+{\bf A}\boldsymbol{\xi}+{\bf B}(Y_{0},Y_{0})^{\top})-(\boldsymbol{\tau}+{\bf A}\boldsymbol{\xi}^{(0)}+{\bf B}(Y_{0}^{(0)},Y_{0}^{(0)})^{\top})\|_{p}\right)$
$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+E\left(\|{\bf B}(Y_{0}-Y_{0}^{(0)},Y_{0}-Y_{0}^{(0)})^{\top}\|_{p}\right)$
$=$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+\|{\bf B}\|_{p}E\left(|Y_{0}-Y_{0}^{(0)}|\right).$
Further, we have that $E(|Y_{0}-Y_{0}^{(0)}|)=|\lambda-\lambda^{(0)}|$; for instance, see proof of Thm 1 by Christou and Fokianos (2014). Hence,
$E\left(\|\boldsymbol{\xi}_{1}-\boldsymbol{\xi}_{1}^{(0)}\|_{p}\right)$$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+\|{\bf B}\|_{p}|\lambda-\lambda^{(0)}|$
$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+\|{\bf B}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{1}$
$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}$
$\leq$ $(\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p})\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p},$
where we have used that $\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{1}\leq 2^{1-1/p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}$. Since $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, the GMC is satisfied, which concludes the proof. $\square$
@@
Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a first-order negative binomial tv-DINGARCH process. Under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.1 holds (for $p=1$) and gives us that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ has a stationary and ergodic solution. Moreover, this process is a Markov chain. Therefore, we need to show that Tweedie’s criterion (for instance, see Chapter 15 from Meyn and Tweedie (1993)) is satisfied for an appropriate test function to prove the finiteness of arbitrary moments.
@@
with $E(Y)=\lambda$, which yields that $E(Y^{d})=\lambda^{d}(1+1/\phi)+O(\lambda^{d-1})$ as $\lambda\rightarrow\infty$. Using this result, it follows that
@@
Hence, (A-4) becomes
@@
as $r\rightarrow\infty$, where the last inequality follows from lemma’s assumption implying that $\beta_{1}+\beta_{2}<1$ and $\alpha_{1}+\alpha_{2}<1$. Furthermore, the remaining term converges to 0 as $r\rightarrow\infty$ because $\dfrac{O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}=O(r^{k-1})/O(r^{k})=O(r^{-1})$.
Therefore, there exist real constants $\kappa_{1}\in(0,1)$, $\kappa_{2}>0$, and $L>0$ such that (A-4) is bounded from above as follows:
@@
In other words, Tweedie’s criterion is satisfied and we conclude that the $k$-th moment of $\lambda_{t}$ and $\phi_{t}$ are finite for arbitrary $k\in\mathbb{N}$. Now, using the fact that $Y_{t}|\mathcal{F}_{t-1}\sim\mbox{NB}(\lambda_{t},\phi_{t})$ and Eq. (A-5), we have that the $k$-th conditional moment of $Y_{t}$ given $\mathcal{F}_{t-1}$ is a polynomial on $\lambda_{t}$ and $1/\phi_{t}$; note that $1/\phi_{t}\leq 1/\alpha_{0}$ for all $t$. Since the $l$-th moment of $\lambda_{t}$ is finite for arbitrary $l\in\mathbb{N}$, we obtain that the unconditional $k$-th moment of $Y_{t}$ is finite for all $k\geq 1$ as well. $\square$
@@
Under the conditions of Theorem 3.1
$\sup\limits_{\boldsymbol{\theta}\in\mathcal{B}}\Big{|}\frac{1}{n}\ell(\boldsymbol{\theta})-\frac{1}{n}\tilde{\ell}(\boldsymbol{\theta})\Big{|}\overset{a.s.}{\longrightarrow}0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{as}\leavevmode\nobreak\ n\rightarrow\infty$
following the notation from Subsection 3.1.
Proof.
Let $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$ denote the stationary and ergodic solution of model (3). In addition, let $\tilde{\boldsymbol{\xi}}_{t}=(\tilde{\lambda}_{t},\tilde{\phi}_{t})^{\top}$ the process obtained with some starting value $\tilde{\boldsymbol{\xi}}_{0}=(\tilde{\lambda}_{0},\tilde{\phi}_{0})^{\top}$. Then, arguing as in Christou and Fokianos (2014, Lemma A.2) and using that $\|{\bf A}\|_{1}<1$. we obtain that
$\sup\limits_{\boldsymbol{\theta}\in\mathcal{B}}\|\boldsymbol{\xi}_{t}-\tilde{\boldsymbol{\xi}}_{t}\|_{1}\leq C\|{\bf A}\|^{t}_{1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ t,$
almost surely from the compactness of $\mathcal{B}$, for some constant $C>0$. Now, consider the differences $\ell_{t}(\boldsymbol{\theta})-\tilde{\ell}_{t}(\boldsymbol{\theta})$ following the notation introduced by (5) and (6). By grouping terms we obtain the following inequalities using repeatedly the facts that for any $x,y>0$ it holds that $|\log(x/y)|\leq|x-y|/\min(x,y)$, $\lambda_{t}>\beta_{0}$, $\phi_{t}>\alpha_{0}$ and $\log x<x-1$, for all $x>0$.
$\bigg{|}y_{t}\Bigl{[}\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})\Bigr{]}-y_{t}\Bigr{[}\log\tilde{\lambda}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi}_{t})\Bigl{]}\bigg{|}$ $\leq 2y_{t}\frac{\|\xi_{t}-\tilde{\xi}_{t}\|_{1}}{\beta_{0}},$
$\bigg{|}\phi_{t}\Bigl{[}\log\phi_{t}-\log(\lambda_{t}+\phi_{t})\Bigr{]}-\tilde{\phi}_{t}\Bigr{[}\log\tilde{\phi}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi}_{t})\Bigl{]}\bigg{|}$ $\leq\|\xi_{t}-\tilde{\xi}_{t}\|_{1}\Bigl{\{}\Bigl{[}(\frac{2}{\alpha_{0}}+1)+\phi_{t}+\lambda_{t}+2\|\xi_{t}-\tilde{\xi}_{t}\|_{1}\Bigr{]}\Bigr{\}},$
$\Bigr{|}\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(y_{t}+\tilde{\phi}_{t})\Bigl{|}\leq(\frac{2}{\alpha_{0}}+1)(y_{t}+\phi_{t}+1)\|\xi_{t}-\tilde{\xi}_{t}\|_{1},$
$\Bigr{|}\log\Gamma(\phi_{t})-\log\Gamma(\tilde{\phi}_{t})\Bigl{|}\leq(\frac{2}{\alpha_{0}}+1)(\phi_{t}+1)\|\xi_{t}-\tilde{\xi}_{t}\|_{1}.$
For the last two inequalities, we employ Stirling’s approximation to the Gamma function. In conclusion
$\sup\limits_{\boldsymbol{\theta}\in\mathcal{B}}\Big{|}\frac{1}{n}\ell(\boldsymbol{\theta})-\frac{1}{n}\tilde{\ell}(\boldsymbol{\theta})\Big{|}$$\leq$ $\frac{C}{n}\sum_{t=1}^{n}\|\boldsymbol{A}\|^{t}_{1}W_{t},$
where $W_{t}$ is a positive linear combination of $(Y_{t},\lambda_{t},\phi_{t})$. Theorem 2.2 implies that $E|W_{t}|^{k}<\infty$, for all $k\in\mathbb{N}$. But using Markovs inequality for every $\epsilon>0$ we have that
$\sum_{t=1}^{\infty}P\Big{(}\|\boldsymbol{A}\|^{t}_{1}W_{t}>\epsilon\Big{)}\leq\displaystyle\sum_{t=1}^{\infty}\frac{\|\boldsymbol{A}\|^{kt}_{1}E(W_{t}^{k})}{\epsilon^{k}}<\infty.$
Hence, $\|\boldsymbol{A}\|^{t}_{1}W_{t}\overset{a.s.}{\longrightarrow}0$, as $t\rightarrow\infty$ and therefore the proof is concluded. ∎
A-3 Proof of Lemma 3.1
We have that $|S_{1t}|\leq|y_{t}/\lambda_{t}-1|\leq y_{t}/\beta_{0}+1$. Furthermore, $\partial\lambda_{t}/\partial\beta_{i}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\lambda_{1}$, for $i=0,1,2$. Using these results and under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.2 gives that the score function associated with $\beta_{0},\beta_{1}$ and $\beta_{2}$ is integrable. We now study the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$. Note that $\left|-\dfrac{y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}\right|\leq\dfrac{y_{t}}{\lambda_{t}+\phi_{t}}+1\leq\dfrac{y_{t}}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)+1$, where we have used that fact that $\dfrac{1}{a+b}\leq\dfrac{1}{2}\left(\dfrac{1}{a}+\dfrac{1}{b}\right)$ for $a,b>0$. Also, by using that $\dfrac{1}{2x}<\log x-\Psi(x)<\dfrac{1}{x}$ for $x>0$ (Alzer, 1997), we obtain that $0<\log\phi_{t}-\Psi(\phi_{t})<\phi_{t}^{-1}<\alpha_{0}^{-1}$. Similarly, it follows that $|\Psi(\lambda_{t}+\phi_{t})-\log(\lambda_{t}+\phi_{t})|<(\lambda_{t}+\phi_{t})^{-1}\leq\dfrac{1}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)$. Therefore, $|S_{2t}|$ is bounded above by a linear combination of $y_{t}$. Moreover, $\dfrac{\partial\phi_{t}}{\partial\alpha_{j}}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\phi_{1}$, for $j=0,1,2$. Again, an application of Theorem 2.2 gives that the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ is also integrable.
Lemma A-2.
@@
We will show that the result holds for the third derivative of $\ell(\cdot)$ with respect to $\beta_{1}$ and also $\alpha_{1}$. The remaining cases are similar and therefore are omitted.
@@
The case for $\alpha_{1}$ is more involved. We have that $\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\alpha_{1}^{3}}=\dfrac{\partial^{2}S_{2,t}}{\partial\phi_{t}^{2}}\left(\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right)^{3}$, where
@@
Inequality (15) from Guo and Qi (2013) gives that $|\Psi^{\prime\prime}(x)|<\dfrac{1}{x}+\dfrac{2}{x^{3}}$, for $x>0$. Hence, it follows that $|\Psi^{\prime\prime}(y_{t}+\phi_{t})|<\dfrac{1}{(y_{t}+\phi_{t})^{2}}+\dfrac{2}{(y_{t}+\phi_{t})^{3}}\leq\dfrac{1}{\phi_{t}^{2}}+\dfrac{2}{\phi_{t}^{3}}\leq\dfrac{1}{\alpha_{0}^{2}}+\dfrac{2}{\alpha_{0}^{3}}\leq\dfrac{1}{\alpha_{0,low}^{2}}+\dfrac{2}{\alpha_{0,low}^{3}}$. Similarly, $|\Psi^{\prime\prime}(\phi_{t})|<\dfrac{1}{\alpha_{0,low}^{2}}+\dfrac{2}{\alpha_{0,low}^{3}}$. The modulus of the remaining terms in (A-7) are bounded above (as done for the third derivative involving $\beta_{1}$) by a linear function of $y_{t}$ with positive coefficients do not depend on $\boldsymbol{\theta}$. Moreover, $\left|\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right|\leq\displaystyle\sum_{j=1}^{t-1}\alpha_{2,up}^{j-1}y_{t-j}\equiv\mu_{2t}$.
@@
A-4 Proof of Theorem 3.1
We will show that the Conditions (A1), (A2), and (A3) from Lemma 3.1 by Jensen and Rahbek (2004) (see also Fokianos et al. (2009) and Christou and Fokianos (2014)) are satisfied, which yield the theorem. As argued before for the general case, we use Proposition 3.1 and the Central Limit Theorem for martingale difference to establish the weak convergence involving the score function in (8). The existence and finiteness of the matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ follow from Theorem 2.2 that provides the finiteness of the moments of all orders for the linear NB tv-DINARCH process under the assumption that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$. Therefore, Condition (A1) holds.
Moreover, we obtain from Theorem 2.1 that $\{Y_{t}\}_{t\in\mathbb{Z}}$ is a stationary and ergodic sequence. Then, the Law of Large Numbers for stationary and ergodic sequences gives us that (9) holds true, with the existence and finiteness of the matrix $\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})$ being ensured again by Theorem 2.2.
We prove next that $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ is non-singular. Suppose that it is singular. Then there exist vectors $\nu_{1}$, $\nu_{2}$ both in $\mathbb{R}^{3}$ such that the six-dimensional vector $\nu=(\nu_{1}^{\top},\nu_{2}^{\top})^{\top}$ satisfies $\nu_{1}^{\top}{\partial\lambda_{t}}/{\partial\boldsymbol{\theta}_{1}}+\nu_{2}^{\top}{\partial\phi_{t}}/{\partial\boldsymbol{\theta}_{2}}=0$, a.s for $t\geq 1$. But both ${\partial\lambda_{t}}/{\partial\boldsymbol{\theta}_{1}}$ and ${\partial\phi_{t}}/{\partial\boldsymbol{\theta}_{2}}$ are stationary so this holds for all $t$. However,
$\frac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{1}}$ $=(1,Y_{t-1},\lambda_{t-1})^{\top}+\beta_{2}\frac{\partial\lambda_{t-1}}{\partial\boldsymbol{\theta}_{1}},$
$\frac{\partial\phi_{t}}{\partial\boldsymbol{\theta}_{2}}$ $=(1,Y_{t-1},\phi_{t-1})^{\top}+\alpha_{2}\frac{\partial\phi_{t-1}}{\partial\boldsymbol{\theta}_{2}}.$
The matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ is block-diagonal. Therefore it is singular if and only if $\nu_{1}^{\top}(1,Y_{t-1},\lambda_{t-1})^{\top}+\nu_{2}^{\top}(1,Y_{t-1},\phi_{t-1})^{\top}=0$ a.s for all $t\geq 1$. But this cannot hold because of the constant term 1 and the fact that $Y_{t}$, $\lambda_{t}$ and $\phi_{t}$ are non-zero, for some $t$.
That is, Condition (A2) from Lemma 3.1 by Jensen and Rahbek (2004) is satisfied. Finally, Condition (A3) has been established in Lemma A-2.